In theory, for freeze concentration, water freezes out as "pure ice" leaving all other non-aqueous material...

  1. In theory, for freeze concentration, water freezes out as "pure ice" leaving all other non-aqueous material in solution. However, some solids always seem to be "caught" in the ice crystals when they are removed.

    In a highly fictitious process, a processor wants to concentrate some low molecular weight material present in a particular plant extract. It is believed to have some health benefits due to its bio-activity. Since this material is considered too delicate to concentrate by heat treatment, the processor decides to use freeze concentration to concentrate the solids in solution to about 5% by weight. The remaining concentrated solution will then undergo further processing to produce a final powdered product.

    The processor starts off with 1,225 kg of solution containing 1.5% of the low molecular weight molecules. Following freeze concentration, 322.1 kg of concentrated solution are obtained containing 5.3% of the desired solids by weight.

    It is suspected that some of the solids are being lost in the ice during the concentration process. (11 marks total)

    1. Calculate the weight of solids lost during the freeze concentration process (if any are indeed lost). Express your answer in kg to two decimal places. (3 marks)

    2. Express the weight of solids lost as a percentage of the total ice removed in the process. Give your answer to two decimal places. (2 marks)

    3. Express the weight of solids lost as a percentage of the initial weight of solids entering the process. Give your answer to one decimal place. (2 marks)

    4. Do you think that this process is functioning in a satisfactory manner? The processor is willing to accept small losses of product (i.e., solids) as long as they are less than 0.5% of the total weight of the ice. (2 marks)

    5. What limits the extent of concentration which can be achieved through freeze-concentration processes? Why (or how)? (2 marks).
Answer #1

As per Chegg Q&A Guidelines, I am answering the first four subparts of the question.

a) The original solution has a mass of 1225 kg of which 1.5% comes from the low molecular weight molecules. Hence the mass due to the low molecular weight molecules in the original solution is 0.015*1225 = 18.375 kg.

After freeze concentration, the mass of concentrated solution becomes 322.1 kg. Of this, 5.3% comes from the low molecular weight molecules. Thus the mass of the desired product is 0.053*322.1 = 17.071 kg.

Thus, the loss of product in the process of concentration is 18.375 - 17.071 = 1.304 kg. Rounding this to two decimal places gives 1.30 kg.

b) The remaining solution was ice. Thus, the amount of ice removed = 1225 - (322.1 + 1.30) = 1225 - 323.4 = 901.6 kg.

Thus, the weight of solids lost as a percentage of ice lost = 1.304 * 100 / 901.6 = 0.14%.

c) The initial weight of solids was 18.375 kg. Thus, the weight of solids lost as a percentage of initial weight of solids = 1.30 * 100 / 18.375 = 7.08%.

d) The total weight of ice = weight of solution - weight of solids = 1225 - 18.375 = 1206.625 kg.

The weight of solids lost as a percentage of total ice = 1.30 * 100 / 1206.625 = 0.11%.

Clearly this loss is less than the threshold 0.5% loss allowed. Hence this process is satisfactory.